### v1

@@ D.1 Proof of Theorem 1
For $i\in\{1,...,L\}$, denote the RR-perturbed $A_{i}$ by ${\tilde{A}}_{i}$. By Proposition 1, for each $i$, ${\tilde{A}}_{i}$ satisfies $\epsilon$-edge-DP. Further, by the parallel composition theorem (McSherry, 2009), $({\tilde{A}}_{1},...,{\tilde{A}}_{L})$ also satisfies $\epsilon$-edge-DP. Finally, by the post-processing property of DP (Proposition 1), Algorithm 1 can achieve $\epsilon$-edge-DP. $\blacksquare$
@@ D.2 Proof of Theorem 2
We will make use of the following result to prove, which is modified from Charisopoulos et al. (2021) using the notation of this paper.
@@
$\|\overline{M}_{l}-Q\|_{2}\leq\frac{\delta}{8},\quad{for\;all\;}l=1,...,L.$ (D.2)
@@
${\mathrm{dist}}(\tilde{V},V)\lesssim\frac{1}{\delta^{2}}\max_{l}\|\overline{M}_{l}-Q\|_{2}^{2}+\frac{1}{\delta}\|\frac{1}{L}\sum_{l=1}^{L}\overline{M}_{l}-Q\|_{2}.$
@@
$\|\overline{M}_{l}-Q\|_{2}$ $\leq\|\overline{M}_{l}-Q_{l}\|_{2}+\|Q_{l}-Q\|_{2}$
$\leq\|\overline{M}_{l}-Q_{l}\|_{2}+C^{\prime}\mathcal{H}(n,\rho,L),$
@@
$\|Q_{l}-Q\|_{2}$ $=\left\|\frac{P_{l}^{2}}{n}-\frac{1}{L}\sum_{i=1}^{L}\frac{P_{i}^{2}}{n}\right\|_{2}\leq\frac{1}{n}\|\Delta\|_{2}^{2}\|B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{i=1}^{L}B_{i}\Delta^{2}B_{i}\|_{2}$
$\leq C\frac{1}{K}\|B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{i=1}^{L}B_{i}\Delta^{2}B_{i}\|_{2}\leq C^{\prime}\mathcal{H}(n,\rho,L),$
@@
with $\Delta:={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$, $\bar{\Theta}:=\Theta\Delta^{-1}$, and $\bar{\Theta}^{\intercal}\bar{\Theta}=\mathbb{I}$, and Assumptions 1 and 4. To meet (D.1) and (D.2), we let $\delta\asymp\lambda_{\min}(Q)$ and enforce
$\|\overline{M}_{l}-Q_{l}\|_{2}\leq\frac{\lambda_{\min}(Q)}{16}\quad{and}\quad\mathcal{H}\leq\frac{\lambda_{\min}(Q)}{16C^{\prime}},$
which follows by our condition in Theorem 2. We can thus use the results in Proposition 3 to obtain that,
$\mathrm{dist}(\tilde{V},V)\lesssim{\max_{l}\frac{\|\overline{M}_{l}-Q_{l}\|_{2}^{2}}{\lambda_{\min}^{2}(Q)}}+{\frac{\mathcal{H}^{2}(n,\rho,L)}{\lambda_{\min}^{2}(Q)}}+\frac{\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}}{\lambda_{\min}(Q)}.$
@@ D.3 Proof of Theorem 3
Consequently, $E_{1}$, $E_{2}$ and $E_{3}$ are the main effects and we obtained the results of Theorem 3. $\blacksquare$
D.4 Proof of Theorem B.1
The proof is similar to the that of Theorem 3. We here go through the main steps.
Recalling (A.2) and (11), we can write
$n\overline{M}_{l}-nQ_{l}:=n\overline{M}_{l}-nP_{l}^{2}:=E_{l,0}+E_{l,1}+E_{l,2}+E_{l,3}-\frac{G_{l}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}.$
Next, we bound these error terms, respectively.
For $E_{l,0}:={-P_{l}}\cdot{\operatorname{diag}}(P_{l})-{\operatorname{diag}}(P_{l})\cdot P_{l}+({\operatorname{diag}}(P_{l}))^{2}$, we can easily show that
$\|E_{l,0}\|_{2}\leq cn\rho^{2}.$
For $E_{l,1}:=X_{l}(P_{l}-{\operatorname{diag}}(P_{l}))+(P_{l}-{\operatorname{diag}}(P_{l}))X_{l}$, using Lemma 1 with the same Bernstein tail condition as those in the proof of Theorem 3, we can obtain that with probability larger than $1-n^{-\beta}$ for some $\beta>0$,
$\|E_{l,1}\|_{2}\leq c\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1}n^{3/2}\rho\log^{1/2}(n),$
provided that $\log^{1/2}\,n\leq C(q\rho+1-q^{\prime})^{1/2}n$, which can be implied by $\log n\leq Cn(1-q^{\prime})=C{n}/({1+{\mathrm{e}}^{\epsilon}})$ by noting our condition $\epsilon\leq C{\log}\,n$.
For $E_{l,2}:=X_{l}^{2}-{\operatorname{diag}}(X^{2}_{l})$, using Lemma 2 with the same Bernstein tail condition as the one used in the proof of Theorem 3, we observe that with probability larger than $1-O(n^{-1})$,
$\|E_{l,2}\|_{2}\leq c\max\left\{vn\log(n),\,\sqrt{v}\sqrt{n}\log^{3/2}(n),\,\log^{2}(n)\right\}/(q+q^{\prime}-1)^{2},$ (D.5)
where $v:=q\rho+1-q^{\prime}$. Note that by the condition that $\epsilon\leq C{\log}\,n$, we have ${\log}\,n\leq Cn(1-q^{\prime})\leq Cn(q\rho+1-q^{\prime})$, and thus the first term in (D.5) dominates and
$\|E_{l,2}\|_{2}\leq c\frac{(q\rho+1-q^{\prime})n\log(n)}{(q+q^{\prime}-1)^{2}}.$
For the last two error terms $E_{l,3}-\frac{G_{l}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}$, it is easy to see that
$\|E_{l,3}-\frac{G_{l}(q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\|_{2}\leq c\frac{(1-q^{\prime}+(q+q^{\prime}-1)\rho)^{2}}{(q+q^{\prime}-1)^{2}}\cdot n.$
It is easy to see that $E_{l,0}\leq CE_{l,1}$. Now we show that $E_{3}\leq CE_{2}$, for which we only need to show
$(1-q^{\prime}+(q+q^{\prime}-1)\rho)^{2}\leq C(q\rho+1-q^{\prime}){\log}\,n.$
Actually, we have $(1-q^{\prime})^{2}\leq C\log\,n(1-q^{\prime})\leq C\log\,n(q\rho+1-q^{\prime})$, and $((q+q^{\prime}-1)\rho)^{2}\leq Cq\rho\log\,n\leq C(q\rho+1-q^{\prime}){\log}\,n$ by noting that $q>1/2$ and $q+q^{\prime}-1\leq 1$. As a result, $E_{l,1}$ and $E_{l,2}$ dominate. After using the union bound, we finally have
$\max_{l}\|\overline{M}_{l}-Q_{l}\|_{2}\leq C\max\left\{\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(n)}{q+q^{\prime}-1},\,\frac{(q\rho+1-q^{\prime})\log n}{(q+q^{\prime}-1)^{2}}\right\},$
with probability larger than $1-O(Ln^{-\gamma})$ for some constant $\gamma>0$. The results of Theorem B.1 are arrived. $\blacksquare$
D.5 Proof of Corollary 1
The proof of Corollary 1 follows directly from that of Theorem 2. First, we need to show that $Q$ is of rank $K$ and provide the lower bound of $\lambda_{K}(Q)$. Indeed, define $\Delta={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$ and denote $\bar{\Theta}:=\Theta\Delta^{-1}$, then $\bar{\Theta}$ is orthogonal and we can write $Q$ as
@@ D.5 Proof of Corollary 1
Next, we specify the $\mathcal{I}_{1},\mathcal{I}_{2}$ and $\mathcal{I}_{3}$ and the corresponding condition in Theorem 2 that $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ need to be smaller enough. Recall the bound for the single matrix $\overline{M}_{l}$ in Theorem B.1, we find that when
$\rho\gtrsim\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1}\cdot\sqrt{\frac{\log(n)}{n}},$ (D.7)
the bound in (B.1) reduces to the first term $\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(n)}{q+q^{\prime}-1}$ which is automatically smaller than $\lambda_{K}(Q)$ provided with (D.7). As a result, $\mathcal{I}_{1}$ is specified under (D.7). For $\mathcal{I}_{2}$, we do not specify it but only require the condition that
$\mathcal{H}\lesssim n\rho^{2}.$
For $\mathcal{I}_{3}$, the bound directly follows from Theorem 3.
As a result, the results of Corollary 1 have been arrived. $\blacksquare$
@@ D.6 Proof of Theorem 4
$\leq C\|\tilde{V}-VO\|_{2}^{2}\leq C\mathcal{S}^{2},$
where $\mathcal{S}:=\mathcal{I}_{1}+\mathcal{I}_{2}+\mathcal{I}_{3}$ with $\mathcal{I}_{i}$’s defined in Corollary 1, the first inequality follows because we assume that ${\tilde{V}^{\prime}}$ is the global solution minimum of the following $k$-means objective and $VO$ is a feasible solution by Lemma 4,
@@
and the last inequality follows from Corollary 1 and the fact that $\min_{O\in\mathcal{O}_{K}}\|\tilde{V}-VO\|_{2}^{2}\lesssim\mathrm{dist}(\tilde{V},V)$.
@@ Appendix E Auxiliary lemmas
Proof. See Lei and Rinaldo (2015); Chen et al. (2021b), among others. $\blacksquare$

### v2

@@ D.1 Proof of Theorem 1
For $l\in\{1,...,L\}$, denote the RR-perturbed $A_{i}$ by ${\tilde{A}}_{i}$. By Proposition 1, for each $i$, ${\tilde{A}}_{i}$ satisfies $\epsilon$-edge-DP. Further, by the parallel composition theorem (McSherry, 2009), $({\tilde{A}}_{1},...,{\tilde{A}}_{L})$ also satisfies $\epsilon$-edge-DP. Finally, by the post-processing property of DP (Proposition 1), Algorithm 1 can achieve $\epsilon$-edge-DP. $\blacksquare$
@@ D.2 Proof of Theorem 2
We will make use of the following result to prove, which is modified from Charisopoulos et al. (2021) using the notation of this work.
@@
$\|\sum_{l\in S_{i}}\overline{M}_{l}/(L/m)-Q\|_{2}\leq\frac{\delta}{8},\quad{for\;all\;}i=1,...,m.$ (D.2)
@@
${\mathrm{dist}}(\tilde{V},V)\lesssim\frac{1}{\delta^{2}}\max_{i\in[m]}\|\sum_{l\in S_{i}}{\overline{M}_{l}}/{(L/m)}-Q\|_{2}^{2}+\frac{1}{\delta}\|\frac{1}{L}\sum_{l=1}^{L}\overline{M}_{l}-Q\|_{2}.$
@@
$\max_{i\in[m]}\|\sum_{l\in S_{i}}{\overline{M}_{l}}/{L/m}-Q\|_{2}$ $\leq\max_{i\in[m]}\|\sum_{l\in S_{i}}{(\overline{M}_{l}-Q_{l})}/{(L/m)}\|_{2}+\max_{i\in[m]}\|\sum_{l\in S_{i}}{Q_{l}}/{(L/m)}-Q\|_{2}$
$\lesssim\max_{i\in[m]}\|\sum_{l\in S_{i}}{(\overline{M}_{l}-Q_{l})}/{(L/m)}\|_{2}+\mathcal{H}(n,\rho,L),$
@@
$\max_{i\in[m]}\|\sum_{l\in S_{i}}{Q_{l}}/{(L/m)}-Q\|_{2}$ $=\max_{i\in[m]}\left\|\frac{1}{L/m}\sum_{l\in S_{i}}\frac{P_{l}^{2}}{n}-\frac{1}{L}\sum_{l=1}^{L}\frac{P_{l}^{2}}{n}\right\|_{2}$
$\leq\max_{i\in[m]}\frac{1}{n}\|\Delta\|_{2}^{2}\left\|\frac{1}{L/m}\sum_{l\in S_{i}}B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{l\in[L]}B_{l}\Delta^{2}B_{l}\right\|_{2}$
$\lesssim\max_{i\in[m]}\frac{1}{K}\left\|\frac{1}{L/m}\sum_{l\in S_{i}}B_{l}\Delta^{2}B_{l}-\frac{1}{L}\sum_{l\in[L]}B_{l}\Delta^{2}B_{l}\right\|_{2}\asymp\mathcal{H}(n,\rho,L),$
@@
with $\Delta:={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$, $\bar{\Theta}:=\Theta\Delta^{-1}$, and $\bar{\Theta}^{\intercal}\bar{\Theta}=\mathbb{I}$, and Assumption 1 and Definition 2. To meet (D.1) and (D.2), we let $\delta\asymp\lambda_{\min}(Q)$ and enforce
$\max_{i\in[m]}\|\sum_{l\in S_{i}}(\overline{M}_{l}-Q_{l})/{(L/m)}\|_{2}\lesssim\frac{\lambda_{\min}(Q)}{16}\quad{and}\quad\mathcal{H}\lesssim\frac{\lambda_{\min}(Q)}{16},$
which follows by our condition in Theorem 2 that $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are smaller than a small constant $c$. We can thus use the results in Proposition 3 to obtain that,
$\mathrm{dist}(\tilde{V},V)\lesssim{\max_{l}\frac{\|\sum_{l\in S_{i}}(\overline{M}_{l}-Q_{l})/{(L/m)}\|_{2}^{2}}{\lambda_{\min}^{2}(Q)}}+{\frac{\mathcal{H}^{2}(n,\rho,L)}{\lambda_{\min}^{2}(Q)}}+\frac{\|\sum_{l}\overline{M}_{l}/L-Q\|_{2}}{\lambda_{\min}(Q)}.$
@@ D.3 Proof of Theorem 3
Consequently, $E_{1}$, $E_{2}$ and $E_{3}$ are the main effects. Noting (12), we obtain the results of Theorem 3. $\blacksquare$
D.4 Proof of Theorem 4
The proof of Theorem 4 follows from that of Theorem 2. First, we need to show that $Q$ is of rank $K$ and provide the lower bound of $\lambda_{K}(Q)$. Indeed, define $\Delta={\operatorname{diag}}(\sqrt{n_{1}},...,\sqrt{n_{K}})$ and denote $\bar{\Theta}:=\Theta\Delta^{-1}$, then $\bar{\Theta}$ is orthogonal and we can write $Q$ as
@@ D.4 Proof of Theorem 4
Next, we specify the upper bounds for $\mathcal{I}_{1},\mathcal{I}_{2}$ and $\mathcal{I}_{3}$ and the corresponding condition in Theorem 2 that $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ need to be smaller enough. Similar to the proof of Theorem 3, we can show that with probability larger than $1-O(m(s+n)^{-\nu})$ for some positive constant $\nu$,
$\small\max_{i}\|\sum_{l\in S_{i}}(\overline{M}_{l}-Q_{l})/|S_{i}|\|_{2}$ $\leq C\max\bigg{\{}\frac{(q\rho+1-q^{\prime})^{1/2}n^{1/2}\rho\log^{1/2}(s+n)}{(q+q^{\prime}-1)\sqrt{s}},$
$\quad\quad\frac{(q\rho+1-q^{\prime})\log(s+n)}{(q+q^{\prime}-1)^{2}\sqrt{s}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}}\bigg{\}}.$ (D.6)
When
$\rho\lesssim\sqrt{\frac{\log(s+n)}{n}}\cdot\frac{(q\rho+1-q^{\prime})^{1/2}}{q+q^{\prime}-1},$ (D.7)
the second term in the RHS of (D.4) dominates the first one. As a result,
$\mathcal{I}_{1}$ $:=\frac{\max_{i}\|\sum_{l\in S_{i}}\overline{M}_{l}/|S_{i}|-Q\|_{2}^{2}}{\lambda_{\min}^{2}(Q)}$
$\lesssim\max\bigg{\{}\frac{(q\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{4}}\cdot\frac{\log^{2}(s+n)}{n^{2}\rho^{4}s},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{4}}{(q+q^{\prime}-1)^{4}n^{2}\rho^{4}}\bigg{\}}:=\mathcal{E}_{1}.$
It is then easy to see that
$n\sqrt{s}\cdot\frac{\rho^{2}(q+q^{\prime}-1)^{2}}{q\rho+1-q^{\prime}}\gtrsim{{\log(s+n)}}$
and
$\frac{(q+q^{\prime}-1)\rho+1-q^{\prime}}{\rho(q+q^{\prime}-1)}\lesssim n^{1/2},$
are sufficient conditions for $\mathcal{I}_{1}\lesssim 1$. For $\mathcal{I}_{2}:=\frac{\mathcal{H}^{2}(n,\rho,L)}{n^{2}\rho^{4}}$, it can be implied by Assumptions 1, 2 and Definition 2 that, $\mathcal{H}\lesssim n\rho^{2}$, and thus $\mathcal{I}_{2}=\mathcal{E}_{2}\lesssim 1$ naturally. For $\mathcal{I}_{3}$, we have by Theorem 3 and (D.7) that,
$\small\mathcal{I}_{3}\lesssim\max\bigg{\{}\frac{(q\rho+1-q^{\prime})\log(L+n)}{(q+q^{\prime}-1)^{2}\sqrt{L}n\rho^{2}},\,\frac{((q+q^{\prime}-1)\rho+1-q^{\prime})^{2}}{(q+q^{\prime}-1)^{2}n\rho^{2}}\bigg{\}}:=\mathcal{E}_{3}.$
Finally, noting (12), we obtain the results of Theorem 4. $\blacksquare$
D.5 Proof of Corollary 1
Following Theorem 4, we only need to compare $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ with $\mathcal{E}_{3}$. When
$m\lesssim\frac{n\rho^{2}(q+q^{\prime}-1)^{2}}{(q\rho+1-q^{\prime})\log(s+n)}\cdot L^{1/2},$
the first term in $\mathcal{E}_{3}$ dominates the first term in $\mathcal{E}_{1}$. The second term in $\mathcal{E}_{3}$ dominates the second term in $\mathcal{E}_{1}$ because the former is just the squared version of the latter and they are both of smaller order than 1. And when
$\mathcal{H}\lesssim\max\bigg{\{}\frac{n^{1/2}\rho\,(q\rho+1-q^{\prime})^{1/2}\log^{1/2}(L+n)}{(q+q^{\prime}-1)L^{1/4}},\,\frac{n^{1/2}\rho\,((q+q^{\prime}-1)\rho+1-q^{\prime})}{q+q^{\prime}-1}\bigg{\}},$
$\mathcal{E}_{3}$ dominates $\mathcal{E}_{2}$. Hence we obtain the results of Theorem 4. $\blacksquare$
@@ D.6 Proof of Theorem 5
$\leq C\|\tilde{V}-VO\|_{2}^{2}\leq C\mathcal{E}_{3}^{2},$
where $\mathcal{E}_{3}$ is defined in Theorem 4, the first inequality follows because we assume that ${\tilde{V}^{\prime}}$ is the global solution minimum of the following $k$-means objective and $VO$ is a feasible solution by Lemma 4,
@@
and the last inequality follows from Theorem 4 and the fact that $\min_{O\in\mathcal{O}_{K}}\|\tilde{V}-VO\|_{2}^{2}\lesssim\mathrm{dist}(\tilde{V},V)$.
@@ Appendix E Auxiliary lemmas
Proof. See Lei and Rinaldo (2015); Chen et al. (2021), among others. $\blacksquare$